# Gate teleportation-assisted routing for quantum algorithms

### v1

@@ Abstract
The limited qubit connectivity of quantum processors poses a significant challenge in deploying practical algorithms and logical gates, necessitating efficient qubit mapping and routing strategies. When implementing a gate that requires additional connectivity beyond the native connectivity, the qubit state must be moved to a nearby connected qubit to execute the desired gate locally. This is typically achieved using a series of SWAP gates creating a SWAP path. However, routing methods relying on SWAP gates often lead to increased circuit depth and gate count, motivating the need for alternative approaches. This work explores the potential of teleported gates to improve qubit routing efficiency, focusing on implementation within specific hardware topologies and benchmark quantum algorithms. We propose a routing method that is assisted by gate teleportation. It establishes additional connectivity using gate teleportation paths through available unused qubits, termed auxiliary qubits, within the topology. To optimize this approach, we have developed an algorithm to identify the best gate teleportation connections, considering their potential to reduce the depth of the circuit and address possible errors that may arise from the teleportation paths. Finally, we demonstrate depth reduction with gate teleportation-assisted routing in various benchmark algorithms, including case studies on the compilation of the Deutsch-Jozsa algorithm and the Quantum Approximation Optimization Algorithm (QAOA) for heavy-hexagon topology used in IBM 127-qubit Eagle r3 processors. Our benchmark results show a $10-20\%$ depth reduction in the routing of selected algorithms compared to regular routing without using the teleported gate.
@@ 1 Introduction
Most of the current routing methods combine mapping and SWAP-based routing schemes. These methods use classical optimization algorithms to map logical qubits of quantum algorithms to physical qubits on quantum hardware and to determine the optimal placement of SWAP gates for routing. These problems can be formulated using graph theory: the mapping problem resembles subgraph isomorphism, while the routing problem is closely related to the token-swapping problem [6]. As the token-swapping problem is classified as NP-hard, the routing problem is thought to belong to the same complexity class [6, 7]. Current routing algorithms are based on various classical techniques, including shortest-path search [8, 9], heuristic methods [10, 11, 12, 13, 14], or other optimization techniques [15, 16, 17, 18, 19, 20, 21, 22, 23]. However, these algorithms suffer from scalability issues and generally produce a suboptimal solution with high SWAP overhead. Recently developed reinforcement learning-based routing methods have demonstrated better scalability and the potential to improve routing efficiency [24, 25]. Software development frameworks such as Qiskit use the SWAP-based bidirectional heuristic search algorithm (SABRE) [10] and its variants. The $\ket{\rm{ket}}$ framework [26] also employs heuristic methods and iterative graph-based approaches to improve performance and ensure compatibility with various quantum hardware architectures.
Quantum state teleportation [27] is a protocol in quantum information theory without a classical analogue that allows one to move a quantum state between two parties over an arbitrary distance using entangled states and classical communication. Similarly, quantum gate teleportation [28] is a protocol that implements quantum gates between spatially separated qubits using shared entanglement in auxiliary qubits and classical communication. Quantum state and gate teleportation have been shown to have applications for quantum communication tasks, such as quantum networks [29], quantum repeaters [30], and quantum cryptography [31]. In quantum computing, quantum teleportation is necessary for measurement-based quantum computing [32], and also serves as primitive for universal fault-tolerant quantum computing [28].
Quantum gate teleportation can be used to implement a gate between the qubits that are not directly connected over arbitrary distance within a quantum processor topology as long as it is possible to connect the starting position to the final position using entangled states through auxiliary qubits and perform classically controlled operations upon measuring those auxiliary qubits.
@@
The implementation of the gate teleportation scheme on a given hardware involves establishing a virtual connection between the control and target qubits using auxiliary qubits. These auxiliary qubits can be selected from the unused qubits in the hardware after mapping all data qubits. For example, if we want to run a 20-qubit algorithm on a 50-qubit device, we will have 30 auxiliary qubits available to establish the teleportation path. Teleporting all non-local gates may not be optimal, as it can increase circuit complexity and accumulate errors. Therefore, a key challenge is to identify which gates should be teleported to achieve optimal routing. We address this critical question of identifying which gates need to be teleported and determining where teleportation gates should be introduced. To this end, we have developed the Routing with Teleported Gates (RTG) method, a systematic approach designed to identify optimal virtual connections and routing pathways utilizing these connections. We introduce gate teleportation as an add-on to existing SWAP-based routing and mapping methods. This hybrid approach selectively employs teleportation for specific gates, balancing the trade-off between circuit depth, gate count, and error, thereby enhancing the overall efficiency of quantum circuit execution. We quantify the results by examining the depth and number of two-qubit gates required for various algorithms. Finally, we present a case study of the routing of the Deutsch-Jozsa and the Quantum Approximation Optimization Algorithm (QAOA) on the heavy-hexagon topology of IBM 127-qubit Eagle r3 processors.
@@ 2 Methods
The teleportation of controlled gates requires both the target and control qubits to be locally connected with pre-entangled auxiliary qubits in a 1D chain topology or at least connected to their nearest neighbours. Teleportation protocol consists of a series of measurements on the auxiliary qubits, followed by conditional operations applied to the target and control qubits based on the outcomes of these measurements. Teleporting a CNOT gate through an auxiliary path with $N$ qubits requires $N+1$ CNOT gate operations and $N$ mid-circuit measurements. However, only two of these CNOT gates operate directly on the data qubits—one acts on the control qubit, and the other acts on the target qubit. Similarly, controlled $U$ teleportation requires $N$ CNOT gate operations, one controlled $U$ gate, and $N$ mid-circuit measurements. Importantly, the depth of these teleportation circuits remains constant regardless of the distance $N$. The circuits for the CNOT and the controlled $U$ teleportations over arbitrary distances are illustrated in Figs. 1 and 2.
@@
The RTG searches through the available virtual connections to select the best ones that can reduce temporal depth compared to standard transpilation without virtual edges. While it is theoretically possible to perform an exact or exhaustive search to evaluate all possible virtual connections, this approach quickly becomes impractical as the number of connections increases. The computational resources required to explore every potential connection grow exponentially with the number of connections and the size of the circuit, making it infeasible for circuits with many auxiliary qubits. Therefore, we develop and benchmark a heuristic search algorithm, which takes into account the connectivity requirements of the circuit. This involves limiting the connections to those that align with the additional connectivity needed for the algorithm (more accurately, the circuit that corresponds to the algorithm) or focusing on nearby connections that satisfy the circuit requirements. Another condition considered is that multiple implementations of the teleported gate may not always yield better performance, given the increased error rates and the cost of reset operations required on the auxiliary qubits after every teleported gate operation through the same path. As a result, the search can be constrained only to consider connections that are not repeatedly used. For example, a limit can be set such that only connections repeated in the whole algorithm two times or fewer are considered. Note that this choice is entirely dependent on the algorithm. Algorithms that result in significantly larger depth circuits may benefit from imposing restrictions to limit consecutive gates routed through the same teleportation paths.
@@
where $P_{\rm g}^{\rm 2q}$ and $P_{\rm g}^{\rm tele}$ are the average error probabilities for native two-qubit gate and teleported two-qubit gates, respectively. The number of native two-qubit gates is denoted by $N_{\rm g}$ and the number of teleported two-qubit gates by $N_{\rm tele}$. The two-qubit error cost estimates the cumulative probability of errors by considering the likelihood of errors during both native and teleported two-qubit operations. The noise-aware RTG method selects connections with lower temporal depth and reduced error costs compared to standard transpilation without virtual edges. It is important to note that a connection is chosen only if both conditions are satisfied. However, it is also possible to assign different weights to these criteria, allowing for a more flexible optimization that balances depth reduction and error budget based on specific hardware characteristics.
At the implementation level, the circuit is first transpiled with the optimal virtual connections identified by the RTG method (step 8. of Fig. 3). These virtual connections are then replaced with the corresponding physical implementation teleportation circuits, enabling efficient routing while adhering to hardware constraints.
@@ 3.1 Benchmark with IBM 127-qubit Eagle r3 processor topology
In Fig. 5, we present the depth reduction achieved using the RTG and noise-aware RTG methods compared to the initial line map topology with SABRE routing. For this analysis, we consider the CNOT gate teleportation and assume the gate time for teleported CNOT gates is $t_{\rm g}^{\rm tele}=3t_{\rm g}^{\rm 2q}$ and average error probability $P_{\rm g}^{\rm tele}=10P_{\rm g}^{\rm 2q}$. It is important to note that current devices exhibit significantly higher error probabilities; for instance, as reported in Ref. [33], $P_{\rm g}^{\rm tele}\approx 100P_{\rm g}^{\rm 2q}$, for a distance of teleportation over 10 qubits. As these values are obtained from the first generation of processors capable for mid-circuit measurements and feedfoward, we chose to use values that we expect to be achievable in future devices through improvements in feed-forward times and gate fidelities. We consider the following 15-qubit benchmark circuits for amplitude estimation: Deutsch-Jozsa algorithm, graph state preparation, quantum approximation optimization algorithm (QAOA), quantum Fourier transform (QFT), and QFT applied to qubits (QFT entangled). The results demonstrate a reduction in circuit depth across all cases except for QAOA. The Deutsch-Jozsa algorithm achieves approximately $18\%$ reduction in depth when utilizing both RTG and noise-aware RTG methods. However, performance decreases for the graph state preparation and QFT benchmarks when noise-aware routing is considered (noise-aware RTG). In the case of QAOA, the benchmark circuit requires multiple parametric two-qubit $Z\otimes Z$ interactions (i.e., $\rm{RZZGate}$), which are then decomposed into two CNOT gates separated by a single qubit Z rotation. As a result, each connection involves more than two CNOT gates, and the heuristic search is constrained to avoid repeated use of the same connections, considering the error of repeated teleportation through the same path. This suggests that teleporting RZZ gates may be a more effective option than teleporting CNOT gates directly in this case. This is indeed the case, see Sec. 3.3, where we demonstrate that the teleportation of controlled-U gates yields better results in the context of QAOA. These findings highlight that the effectiveness of the RTG method is highly dependent on the structure of an algorithm and the specific gate operations involved.
@@ 3.2 Case study 1: Deutsch-Jozsa algorithm with CNOT gate teleportation
In Sec. 3.1, we have demonstrated depth reduction by incorporating optimal virtual edges into the coupling map. However, we have not yet included the actual teleportation circuits for these connections (step 8. of Fig. 3). Therefore, the observed depth reduction represents the maximum possible improvement one can expect with virtual edges alone. Here, we present the complete routing process, including the full teleportation circuits utilizing auxiliary qubits, for the Deutsch-Jozsa algorithm on the IBM topology. The benchmark circuit from Ref. [37] is utilized to evaluate the impact of virtual edges on transpilation outcomes at the implementation level. As a baseline, the circuit is first transpiled with 999915151515 qubits arranged in a line topology. After this, we incorporate the optimal virtual edges identified by the RTG method for CNOT gate teleportation into the circuit. We then compare the depth and number of CNOT gates in the baseline line topology with those in the physically routed implementation that includes the virtual edges. To ensure a fair comparison, we also consider the number of CNOT gates executed between data qubits in the teleportation-based implementation, as well as the number of teleported CNOT gates. The results are presented in Table 1.
@@
The results demonstrate a reduction in circuit depth compared to the corresponding line coupling map; however, this reduction is less than anticipated with an ideal coupling map incorporating virtual connections (Virtual c. map). This discrepancy occurs because the ideal coupling map regards virtual connections as equivalent to native connections without accounting for physical implementation, leading to an overestimation of depth reduction. In the case of physical implementation, the total number of CNOT gates is considerably higher than in the line topology. Nonetheless, the number of CNOT gates directly acting on data qubits remains relatively similar. The additional increase is attributed to the auxiliary qubits used for teleportation, which necessitate extra CNOT gates to establish the teleportation paths, see Fig. 1.

### v2

@@ Abstract
The limited qubit connectivity of quantum processors poses a significant challenge in deploying practical algorithms and logical gates, necessitating efficient qubit mapping and routing strategies. When implementing a gate that requires additional connectivity beyond the native connectivity, the qubit state must be moved to a nearby connected qubit to execute the desired gate locally. This is typically achieved using a series of SWAP gates creating a SWAP path. However, routing methods relying on SWAP gates often lead to increased circuit depth and gate count, motivating the need for alternative approaches. This work explores the potential of teleported gates to improve qubit routing efficiency, focusing on implementation within specific hardware topologies and benchmark quantum algorithms. We propose a routing method that is assisted by gate teleportation. It establishes additional connectivity using gate teleportation paths through available unused qubits, termed auxiliary qubits, within the topology. To optimize this approach, we have developed an algorithm to identify the best gate teleportation connections, considering their potential to reduce the depth of the circuit and address possible errors that may arise from the teleportation paths. Finally, we demonstrate depth reduction with gate teleportation-assisted routing in various benchmark algorithms, including case studies on the compilation of the Deutsch-Jozsa algorithm and the Quantum Approximation Optimization Algorithm (QAOA) for heavy-hexagon topology used in IBM 127-qubit Eagle r3 processors. Our benchmark results show a $10-25\%$ depth reduction in the routing of selected algorithms compared to regular routing without using teleported gates.
@@ 1 Introduction
Most of the current routing methods combine mapping and SWAP-based routing schemes. These methods use classical optimization algorithms to map logical qubits of quantum algorithms to physical qubits on quantum hardware and to determine the optimal placement of SWAP gates for routing. These problems can be formulated using graph theory: the mapping problem resembles subgraph isomorphism, while the routing problem is closely related to the token-swapping problem [6]. As the token-swapping problem is classified as NP-hard, the routing problem is believed to belong to the same complexity class [6, 7]. Current routing algorithms are based on various classical techniques, including shortest-path search [8, 9], heuristic methods [10, 11, 12, 13, 14], or other optimization techniques [15, 16, 17, 18, 19, 20, 21, 22, 23]. However, these algorithms suffer from scalability issues and generally produce a suboptimal solution with high SWAP overhead. Recently developed reinforcement learning-based routing methods have demonstrated better scalability and the potential to improve routing efficiency [24, 25]. Software development frameworks such as Qiskit use the SWAP-based bidirectional heuristic search algorithm (SABRE) [10] and its variants. The $\ket{\rm{ket}}$ framework [26] also employs heuristic methods and iterative graph-based approaches to improve performance and ensure compatibility with various quantum hardware architectures.
Quantum state teleportation [27] is a protocol in quantum information theory without a classical analogue that allows one to move a quantum state between two parties over an arbitrary distance using entangled states and classical communication. Similarly, quantum gate teleportation [28] is a protocol that implements quantum gates between spatially separated qubits using shared entanglement in auxiliary qubits and classical communication. Quantum state and gate teleportation have been shown to have applications for quantum communication tasks, such as quantum networks [29], quantum repeaters [30], and quantum cryptography [31]. In quantum computing, quantum teleportation is necessary for measurement-based quantum computing [32], and also serves as a primitive for universal fault-tolerant quantum computing [28].
@@
Qubit routing methods typically move around the state of a qubit using SWAP operations. Similarly, quantum teleportation can be used to teleport the state of a qubit to an arbitrary position within the topology of the quantum computer, as long as it is possible to connect the starting position to the final position using entangled states through auxiliary qubits and perform classically controlled operations upon measuring those auxiliary qubits. An in-depth analysis of the speed-up of the local operations and classical communication (LOCC) assisted routing technique over SWAP-based methods is studied in Ref. [35], as long as $\mathcal{O}(n)$ auxiliary qubits are allowed. In Ref. [36], quantum teleportation is proposed as a complementary method in routing, where the possible teleportation channels through ancillary qubits add virtual edges to the coupling map of a given quantum hardware topology, increasing the possible paths available to the A-based mapping algorithm [37], thus achieving a lower circuit depth. In Ref. [38], a quantum routing method with state teleportation is proposed using entanglement generating devices such as photonic interconnects, creating the required channel for teleportation by entangling qubits with photons, and then performing Bell measurements to swap the entanglement from a pair of qubit-photon entangled states to a qubit-qubit EPR pair, without the need for additional ancillary qubits and gates within the QPU.
In a similar way to quantum state teleportation, quantum gate teleportation can be used to implement a gate between the qubits that are not directly connected over arbitrary distances within a quantum processor topology. The implementation of the gate teleportation scheme on a given hardware involves establishing a virtual connection between the control and target qubits using auxiliary qubits. These auxiliary qubits can be selected from the unused qubits in the hardware after mapping all data qubits. For example, if we want to run a 20-qubit algorithm on a 50-qubit device, we will have 30 auxiliary qubits available to establish the teleportation path. Teleporting all non-local gates may not be optimal, as it can increase circuit complexity and accumulate errors. Therefore, a key challenge is to identify which gates should be teleported to achieve optimal routing. We address this critical question of identifying which gates need to be teleported and determining where teleportation gates should be introduced. To this end, we have developed the Routing with Teleported Gates (RTG) method, a systematic approach designed to identify optimal virtual connections and routing pathways utilizing these connections. We introduce gate teleportation as an add-on to existing SWAP-based routing and mapping methods. This hybrid approach selectively employs teleportation for specific gates, balancing the trade-off between circuit depth, gate count, and error, thereby enhancing the overall efficiency of quantum circuit execution. We quantify the results by examining the depth and number of two-qubit gates required for various algorithms. Finally, we present a case study of the routing of the Deutsch-Jozsa and the Quantum Approximation Optimization Algorithm (QAOA) on the heavy-hexagon topology of IBM 127-qubit Eagle r3 processors.
@@ 2 Methods
The teleportation of controlled gates requires both the target and control qubits to be locally connected with pre-entangled auxiliary qubits in a 1D chain topology or at least connected to their nearest neighbours. Teleportation protocol consists of a series of measurements on the auxiliary qubits, followed by conditional operations applied to the target and control qubits based on the outcomes of these measurements. Teleporting a CNOT gate through an auxiliary path with $N$ qubits requires $N+1$ CNOT gate operations and $N$ mid-circuit measurements. However, only two of these CNOT gates operate directly on the data qubits—one acts on the control qubit and the other acts on the target qubit. Similarly, controlled $U$ teleportation requires $N$ CNOT gate operations, one controlled $U$ gate, and $N$ mid-circuit measurements. Importantly, the depth of these teleportation circuits remains constant regardless of the distance $N$. The circuits for the CNOT and the controlled $U$ teleportations over arbitrary distances are illustrated in Figs. 1 and 2.
@@
The RTG searches through the available virtual connections to select the best ones that can reduce temporal depth compared to standard transpilation without virtual edges. While it is theoretically possible to perform an exact or exhaustive search to evaluate all possible virtual connections, this approach quickly becomes impractical as the number of connections increases. The computational resources required to explore every potential connection grow exponentially with the number of connections and the size of the circuit, making it infeasible for circuits with many auxiliary qubits. Therefore, we develop and benchmark a heuristic search algorithm, which takes into account the connectivity requirements of the circuit. The algorithm primarily employs a greedy heuristic combined with randomized transpilation with the SABRE routing method. It begins by preprocessing the circuit (more accurately, the circuit that corresponds to the algorithm) to identify pairs of qubits that require connectivity beyond what is available in the initial layout, using a directed acyclic graph (DAG) representation. From the set of all available virtual connections, the algorithm selects a subset that is either directly required or spatially close to the required qubit pairs (step 4 in Fig. 3). Another condition considered is that multiple implementations of the teleported gate may not always yield better performance, given the increased error rates and the cost of reset operations required on the auxiliary qubits after every teleported gate operation through the same path. As a result, the search can be constrained to only consider connections that are not used repeatedly. For example, a limit can be set such that only connections repeated twice or less in the whole algorithm are considered. Following this, the algorithm generates all valid combinations of subsets of virtual connections, ensuring that no two connections in a subset share the same qubits. This pruning strategy helps filter out combinations that reuse the same set of qubits for teleportation. For each subset of virtual connections, the algorithm computes the cost (temporal depth) and selects the one with the lowest value. The subset that minimizes the cost is stored as the best solution. Multiple transpilation trials are performed for each subset using different random seeds to find the lowest cost. The randomized search approach mitigates the risk of local minima. At the implementation level, the circuit is first transpiled with the best virtual connections identified by the RTG method (step 8. of Fig. 3). These virtual connections are then replaced with the corresponding physical implementation teleportation circuits, enabling efficient routing while adhering to hardware constraints.
@@
where $P_{\rm g}^{\rm 2q}$ and $P_{\rm g}^{\rm tele}$ are the average error probabilities for native two-qubit gate and teleported two-qubit gates, respectively. The number of native two-qubit gates is denoted by $N_{\rm g}$ and the number of teleported two-qubit gates by $N_{\rm tele}$. The two-qubit error cost estimates the cumulative probability of errors by considering the likelihood of errors during both native and teleported two-qubit operations. The noise-aware RTG method selects connections with lower temporal depth and reduced error costs compared to standard transpilation without virtual edges. For each subset of virtual connections, the algorithm computes the temporal depth and the error cost and selects the one with the lowest values for both. The subset that minimizes both costs is stored as the best solution. However, it is also possible to assign different weights to these criteria, allowing for a more flexible optimization that balances depth reduction and error budget based on specific hardware characteristics.
@@ 3.1 Benchmark with IBM 127-qubit Eagle r3 processor topology
In Fig. 5, we present the depth reduction achieved using the RTG and noise-aware RTG methods compared to the initial line map topology with SABRE routing. For this analysis, we consider the CNOT gate teleportation and assume the gate time for teleported CNOT gates is $t_{\rm g}^{\rm tele}=3t_{\rm g}^{\rm 2q}$ and average error probability $P_{\rm g}^{\rm tele}=10P_{\rm g}^{\rm 2q}$. It is important to note that current devices exhibit significantly higher error probabilities; for instance, as reported in Ref. [33], $P_{\rm g}^{\rm tele}\approx 100P_{\rm g}^{\rm 2q}$, for a distance of teleportation over 10 qubits. As these values are obtained from the first generation of processors capable for mid-circuit measurements and feed-forward, we chose to use values that we expect to be achievable in future devices through improvements in feed-forward times and gate fidelities. We consider the following 15-qubit benchmark circuits for amplitude estimation: Deutsch-Jozsa algorithm, graph state preparation, quantum approximation optimization algorithm (QAOA), quantum Fourier transform (QFT), and QFT applied to entangled qubits (QFT entangled). The results demonstrate a reduction in circuit depth across all cases except for QAOA. The Deutsch-Jozsa algorithm achieves approximately $18\%$ reduction in depth when utilizing both RTG and noise-aware RTG methods. However, performance decreases for the graph state preparation and QFT benchmarks when noise-aware routing is considered (noise-aware RTG). In the case of QAOA, the benchmark circuit requires multiple parametric two-qubit $Z\otimes Z$ interactions (i.e., $\rm{RZZGate}$), which are then decomposed into two CNOT gates separated by a single qubit Z rotation. As a result, each connection involves more than two CNOT gates, and the heuristic search is constrained to avoid repeated use of the same connections, considering the error of repeated teleportation through the same path. This suggests that teleporting RZZ gates may be a more effective option than teleporting CNOT gates directly in this case. This is indeed the case, see Sec. 3.3, where we demonstrate that the teleportation of controlled-U gates yields better results in the context of QAOA. These findings highlight that the effectiveness of the RTG method is highly dependent on the structure of an algorithm and the specific gate operations involved. The RTG method discussed here focuses mainly on teleporting two-qubit gates. However, the framework can be extended to incorporate other teleported gates, such as Toffoli and fanout gates, or to enable constant-depth execution methods [43, 44] for quantum circuits for specific applications.
@@ 3.2 Case study 1: Deutsch-Jozsa algorithm with CNOT gate teleportation
In Sec. 3.1, we have demonstrated depth reduction by incorporating optimal virtual edges into the coupling map. However, we have not yet included the actual teleportation circuits for these connections (step 8. of Fig. 3). Therefore, the observed depth reduction represents the maximum possible improvement one can expect with virtual edges alone. Here, we present the complete routing process, including the full teleportation circuits utilizing auxiliary qubits, for the Deutsch-Jozsa algorithm on the IBM topology. The benchmark circuit from Ref. [41] is utilized to evaluate the impact of virtual edges on transpilation outcomes at the implementation level. As a baseline, the circuit is first transpiled with 999915151515 qubits arranged in a line topology. After this, we incorporate the optimal virtual edges identified by the RTG method for CNOT gate teleportation into the circuit. We then compare the depth and number of CNOT gates in the baseline line topology with those in the physically routed implementation that includes the virtual edges. The “Virtual c. map” represents the case where possible additional connections are added to the initial coupling map without considering the actual implementation of virtual connections. This situation represents an ideal case, demonstrating the maximum reduction in circuit depth that can be achieved with a specific set of virtual connections. To ensure a fair comparison, we also consider the number of CNOT gates executed between data qubits in the teleportation-based implementation, as well as the number of teleported CNOT gates. The results are presented in Table 1.
@@
The results demonstrate a reduction in circuit depth compared to the corresponding line coupling map; however, this reduction is less than anticipated with an ideal coupling map incorporating virtual connections (Virtual c. map). This discrepancy occurs because the ideal coupling map regards virtual connections as equivalent to native connections without accounting for physical implementation, leading to an overestimation of depth reduction. In the case of physical implementation, the total number of CNOT gates is considerably higher than in the line topology. Nonetheless, the number of CNOT gates directly acting on data qubits remains relatively similar. The additional increase is attributed to the auxiliary qubits used for teleportation, which necessitate extra CNOT gates to establish the teleportation paths, see Fig. 1. Reducing circuit depth and the number of gates on data qubits is crucial for optimizing quantum algorithms on near-term devices, where limited coherence times and gate fidelities constrain performance. A shallower circuit helps to reduce decoherence, resulting in improved overall fidelity. Therefore, the reduction in circuit depth discussed here is highly relevant for near-term quantum devices.